OpenStudy (countrygirl1431):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (anonymous):

So write an equation?

OpenStudy (countrygirl1431):

yes

OpenStudy (anonymous):

(y+4)=3/5(x-2) Your formula is (y-y)=m(x-x) So 3/5 is m which is slope, and (x,y) which is (2,-4) This is how I think you do it.